OpenStudy (anonymous):

what are the intercepts of this equation in help me with showing the work 1.25x+1x(less than equal to sign)15

OpenStudy (saifoo.khan):

\[1.25x + x \le 15\]Solve for x \[2.5x \le 15\]\[x \le \frac{15}{2.5}\]

OpenStudy (saifoo.khan):

\[x \le 6\]

OpenStudy (anonymous):

thank you